# envname-P envname#1

### v1

@@ C.1 Semidiscrete optimal transport
To solve for the semidiscrete dual potential $\mathbf{g}$ in practice, we implement Algorithm 1 using JAX and utilize built-in data parallelism to scale to handle large datasets over multiple GPUs and nodes. Since the problem equation $\mathrm{S}_{\varepsilon}$ is concave (and strictly, for $\varepsilon>0$) one can use a variety of different optimizers and obtain a suitable convergence behavior. We use the dot-product cost and scale $\varepsilon$ with the standard deviation of the cost, as proposed by Zhang et al. (2025). In the class-conditional setting, we choose $\beta=10^{2}$ for ImageNet-32 and $\beta=2\times 10^{2}$ for ImageNet-64 and PetFace.
@@
To estimate $\chi^{2}(\mathbf{m}({\mathbf{g}})\,\|\,\mathbf{b})$, we use a total of $2^{20}$ noise samples, which we divide into batches of size $2^{13}$ that can further be sharded along multiple devices, and average equation 3.5 over these batches. For the ImageNet-32x32 dataset, this calculation takes less than 2 minutes on a node of 8 H100 GPUs.
@@ C.2 Flow model training
Images from CelebA are first rescaled to 64x64 resolution. From each image $\mathbf{x}$, low-resolution, noisy images are constructed by first downscaling to 16x16 (4x SR) or 8x8 (8x SR) and adding Gaussian noise with standard deviation $\sigma=0.1$ or $0.2$, i.e. $\mathbf{z}=\texttt{downscale}(\mathbf{x})+\mathbf{\eta}$, $\mathbf{\eta}=\mathcal{N}(0,\sigma^{2}\operatorname{\mathbf{I}})$. The sampled corrupted image $\mathbf{z}$ is treated as the condition paired with $\mathbf{x}$, and we solve conditional optimal transport with $\beta=25$. Flow matching models are then trained to generate samples of 64x64 images $\mathbf{x}$ conditional on downscaled, noisy observations $\mathbf{z}$. We parameterize the flow with a U-Net with the same hyperparameter choices as for 64x64 ImageNet, and conditions are handled by resizing the low-resolution 8x8 or 16x16 image to 64x64 and stacking together with the input to the U-Net. A batch size of 128 and a learning rate of 0.0002 are used for FM training for a total of 500 000500\,000 iterations.Apple and the Apple logo are trademarks of Apple Inc., registered in the U.S. and other countries and regions.

### v2

@@ C.1 Semidiscrete optimal transport
To solve for the semidiscrete dual potential $\mathbf{g}$ in practice, we implement Algorithm 1 using JAX and utilize built-in data parallelism to scale to handle large datasets over multiple GPUs and nodes. Since the problem ($\mathrm{S}_{\varepsilon}$) is concave (and strictly, for $\varepsilon>0$) one can use a variety of different optimizers and obtain a suitable convergence behavior. We use the dot-product cost and scale $\varepsilon$ with the standard deviation of the cost, as proposed by Zhang et al. (2025). In the class-conditional setting, we choose $\beta=10^{2}$ for ImageNet-32 and $\beta=2\times 10^{2}$ for ImageNet-64 and PetFace.
@@
To estimate $\chi^{2}(\mathbf{m}({\mathbf{g}})\,\|\,\mathbf{b})$, we use a total of $2^{20}$ noise samples, which we divide into batches of size $2^{13}$ that can further be sharded along multiple devices, and average (3.5) over these batches. For the ImageNet-32x32 dataset, this calculation takes less than 2 minutes on a node of 8 H100 GPUs.
@@ C.2 Flow model training
Images from CelebA are first rescaled to 64x64 resolution. From each image $\mathbf{x}$, low-resolution, noisy images are constructed by first downscaling to 16x16 (4x SR) or 8x8 (8x SR) and adding Gaussian noise with standard deviation $\sigma=0.1$ or $0.2$, i.e. $\mathbf{z}=\texttt{downscale}(\mathbf{x})+\mathbf{\eta}$, $\mathbf{\eta}=\mathcal{N}(0,\sigma^{2}\operatorname{\mathbf{I}})$. The sampled corrupted image $\mathbf{z}$ is treated as the condition paired with $\mathbf{x}$, and we solve conditional optimal transport with $\beta=25$. Flow matching models are then trained to generate samples of 64x64 images $\mathbf{x}$ conditional on downscaled, noisy observations $\mathbf{z}$. We parameterize the flow with a U-Net with the same hyperparameter choices as for 64x64 ImageNet, and conditions are handled by resizing the low-resolution 8x8 or 16x16 image to 64x64 and stacking together with the input to the U-Net. A batch size of 128 and a learning rate of 0.0002 are used for FM training for a total of 500 000500\,000 iterations.
@@ Appendix D Additional experimental results
Appendix E Sensitivity of dual potential to $\varepsilon$
In 17, we provide a fine-grained picture of the impact of $\varepsilon$ when computing $\mathbf{g}$ on the same ImgN64 dataset (in full dimensions).
Figure 17: We compare the optimal dual potentials $\mathbf{g}^{\star}_{\varepsilon}$ for 3 different levels of $\varepsilon$ using scatter plots. Each scatter plot depicts 2.56M points, where 2D points $((g^{\star}_{\varepsilon})_{j},(g^{\star}_{\varepsilon^{\prime}})_{j})$ are plotted for different $\varepsilon\neq\varepsilon^{\prime}$. As can be shown, the potential vector $\mathbf{g}^{\star}_{\varepsilon}$ varies smoothly, for each datapoint, with $\varepsilon$. These computations are carried out on full-dimension $d=12,288$ on ImgN64.
Appendix F SD-FM: Ablations on ImgN64
F.1 Distribution of MIPS Sampling as a function of $\tau$
Figure 18 proposes to study the distribution of selected images in the dataset of $N$ ImgN64 (full-dimension) as SD-FM training occurs. This pictures provides a complementary view to Fig. 3 showing why the FID metric improve as the $\chi_{2}$ metric decreases.
Figure 18: Effect of varying threshold $\tau$ (in Alg.1) when computing $\mathbf{g}^{\star}_{\varepsilon}$ on the distribution of selected images in the dataset, for $\varepsilon=0.0$ on the left, $\varepsilon=0.01$ right. As can be seen, a loose threshold of $\tau=1$ (essentially no optimization for $\mathbf{g}$) results in images that are never sampled, whereas a finer threshold of $0.05$ equalizes the distribution of sampled images. These computations are carried out on PCA $k=1000$ of the data space originally in full-dimension $d=12,288$, on ImgN64.
F.2 SD-FM Metrics a function of $\tau,\varepsilon$ and cost ablation
We simultaneously study in the following experiments three different factors that lie at the heart of SD-FM:
  • does it matter that the potential $\mathbf{g}^{\star}_{\varepsilon}$ for a certain $\varepsilon$ is truly optimal, in the sense that it respects marginals and guarantees that every point in the dataset is properly sampled? In other words, does low $\tau$ always correlate with better FID and curvature metrics?
  • does it matter for SD-FM to work to narrow down on the optimal transport as defined using the "traditional" $\ell_{2}$ cost $c(x,y)=-\langle x,y\rangle$? Would results be different if one were to select different costs? To answer this we consider three other costs: the "anti-OT" cost $c(x,y)=\langle x,y\rangle$ that promotes, on the contrary, the longest paths, and two random corruptions of the $c_{\sigma}(x,y)=-\langle\sigma\circ x,y\rangle$. The $\sigma$ holds the realizations of $d$ random (Rademacher) variable in 1,1{-1,1} with probability 30% or 70% of taking the value 1-1. This value is sampled only once per experiment.
  • does it matter to have a deterministic association ($\varepsilon=0$) or is it better to inject more randomness ($\varepsilon>0$), having in mind that infinite $\varepsilon$ is equivalent to IFM.
While all of Figures 19, 20, 21, 22 provide a wide wealth of information, we can confidently say, by looking at all of them, that apart from FID in the lowest NFE-4 case, which is slightly more nuanced, all metrics improve as $\varepsilon$ and $\tau$ decrease (closer to "sharp" OT) and clearly highlight the crucial role of the $\ell_{2}$ cost (i.e. not "any" OT plan would work, that with shortest paths results in better performance).
Figure 19: Low $\tau=0.05$ regime (i.e. close to exact SD-OT computation).
Figure 20: Middle $\tau=0.2$ regime (i.e. exact SD-OT computation).
Figure 21: High $\tau=1.0$ regime (no marginal preservation is guaranteed).
Figure 22: Original cost $c(x,y)=-\langle x,y\rangle$, low $\varepsilon=0$ regularization, ablating $\tau$.
F.3 Training SD-OT on Pixel Space vs. Latent Space
F.4 SD-FM: Ablations on ImgN64
Figure 23: We compare the evolution of the $\chi_{2}$ convergence criterion when computing SD-OT on either ImgN32 in original pixel space ($d=3072$) vs. ImgN256 in latent space ($d=4096$). Here the total dataset size is the same (N=2.56N=2.56M). While one would expect SD-OT to converge faster for a smaller dimensional space (i.e. ImgN32), this is not the case, because of the far more regular distribution of latents that is designed to resemble a Gaussian.